Rochester Heating Systems, Inc. calculates its predetermined overhead rate on a quarterly basis. The following estimates were made for the current year Quarterly Predetermined Overhead Estimated Estimated Rate (per Manufacturing Direct-Labor direct-labor Hours 25,000 16,000 Overhead hour) First quarter 75,000 Second 96,000 quarter Third quarter 37,500 Fourth 12,500 84,000 14,000 quarter Total $292,500 67,500 The firms main product, part number SC71, requires $200 of direct material and 20 hours of direct labor per unit. The labor rate is $15 per hour4. Calculate the companys predetermined overhead rate for the year if the rate is calculated annually.(Round your answer to 2 decimal places.) Predetermined overhead ate per hour
Answer #1

Pre-determined overhead rate= (Budgeted total Manufacturing overhead cost/Budgeted direct labor hours)

Quarter Budgeted Mfg Overhead(A) Budgeted Direct Labor Hours(B) Pre-determined Ohd Rate=(A/B)
First $75000 25000 75000/25000=$3 per DLH
Second $96000 16000 96000/16000=$6 per DLH
Third $37500 12500 37500/12500=$3 per DLH
Forth $84000 14000 84000/14000=$6 per DLH
Total $292500 67500 292500/67500=$4.33 per DLH

So, The pre-determined Ohd rate if rate calculated annually= Total Budgeted Mfg Ohd for the year/Total Budgeted DLH

=292500/67500 = $4.33 per DLH